OpenStudy (anonymous):

Given BE=2x+6 And ED=5x-12 in parallelogram ABCD, Find BD.

OpenStudy (anonymous):

Yes it certainly would.

OpenStudy (anonymous):



OpenStudy (anonymous):

sorry i cant draw good its my first time lol

OpenStudy (anonymous):

Are BD and AC diagonals of the parallelogram? @MathSuckzz

OpenStudy (anonymous):

I think so, it doesn't say

OpenStudy (anonymous):

BD = BE + ED BD = (2x + 6) + (5x - 12) BD = 2x + 6 + 5x - 12 BD = 7x - 6

OpenStudy (anonymous):

Does that make sense to you? @MathSuckzz

OpenStudy (anonymous):

yes it does but is the equation supposed to be the final answer?

OpenStudy (anonymous):

Yes it is. Since we don't have a value for x given, that would be the final answer.

OpenStudy (anonymous):

oh okay, thank you so much!